# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v2

@@ 3.1. Generation Stage
FL enables collaborative model training across different energy producers and consumers, enhancing the accuracy of energy generation forecasts and enabling dynamic resource allocation without the need to centralize sensitive operational data (lee2023towards; gupta2023fedgrid). In the context of energy management (abdellatif2024sdcl), FL supports the development of intelligent systems that can predict and optimize power generation patterns while keeping sensitive user data local on smart devices. This approach improves grid efficiency and facilitates more effective demand response programs.
Energy Forecasting In the work by (lee2023towards), an advanced algorithm for FL is proposed to tackle the challenges of non-IID data in distributed solar energy grids. This framework is specifically designed to handle the heterogeneous nature of the data across different solar energy producers, enhancing the model’s ability to converge efficiently in decentralized settings. By addressing the variations in data distribution, the framework improves the accuracy and reliability of predictions, leading to better performance in solar energy generation and more effective resource allocation.
Energy Optimization The FedGrid framework introduced by (gupta2023fedgrid) offers a secure FL solution for optimizing energy management within smart grids. This framework emphasizes the importance of protecting privacy and enhancing data security during the optimization process. By enabling efficient collaboration between various grid components without compromising data confidentiality, FedGrid helps optimize energy distribution and usage. The approach also addresses the key challenges associated with data sharing and collaboration in smart grids, ultimately improving the overall efficiency and reliability of the system. As shown in Figure 3, energy optimization is crucial for coordinating multiple energy resources and consumption, especially when using FL techniques to protect data privacy.
Energy Management In (abdellatif2024sdcl), a Secure, Distributed, and Collaborative Learning (SDCL) framework is presented to enhance the operations of smart grids. This framework focuses on maintaining data privacy and security while facilitating collaborative learning among grid components. By incorporating advanced AI and blockchain technologies, SDCL aims to optimize grid performance, improve system efficiency, and ensure reliable energy management. Additionally, (lin2024privacy) introduces a privacy-preserving energy management framework that utilizes vertical FL for smart data cleaning in smart homes. This approach enhances user privacy while enabling effective analysis and management of smart meter data, ensuring that energy usage is optimized without compromising sensitive information.
@@ 3.2. Transmission and Distribution Stages
For energy distribution (mourtzis2023integration), FL enables optimal energy distribution solutions and flexibility predictions, significantly improving the efficiency of power distribution in SGs. In the area of attack (tran2023efficient; zhao2024federated) and anomaly detection (kesici2024detection; jithish2023distributed), FL has demonstrated its effectiveness in identifying cyber attacks and anomalies within SG systems. This approach allows multiple stakeholders to collaboratively train robust detection models while maintaining data privacy, as raw consumption data remains local. In the realm of assessment (li2024privacy; ren2023qfdsa; ren2023efeddsa), FL has been applied to dynamic security assessment, enabling participants to collaboratively build models while preserving data privacy.
Energy Distribution Mourtzis et al. (mourtzis2023integration) develop an integrated FL framework aims at enhancing the efficiency and security of energy distribution in smart grids. This framework addresses privacy concerns by enabling collaborative learning without sharing raw data, which is particularly relevant in smart grid data analysis. Specifically, they propose an incentive mechanism for FL, edge devices are compensated by service providers based on their contribution to a given task. The payoff function for an edge device $i$ participating in task $\Upsilon_{j,k}$ under service provider $j$ is:
(4) $u_{i,j,k}(Q_{i,j,k},S_{i,j,k})=\varpi\kappa_{p}p_{i,j,k}\Psi_{i,j,k}-(1-\varpi)C_{i,j,k},$
where $\varpi$ is a weight parameter that balances revenue and cost, $\kappa_{p}$ is an adjustment parameter, $p_{i,j,k}$ is the payment to edge device $i$, and $\Psi_{i,j,k}$ is the highest Age of Local Model achieved by $i$ in task $\Upsilon_{j,k}$. The total cost $C_{i,j,k}$ for edge device $i$ includes local data cost, computational cost, and communication cost. The payoff function for service provider $j$ in task $k$ is given by:
(5) $\pi_{j,k}=\mu\nu_{a}\Lambda_{j,k}-(1-\mu)\left(\sum_{i\in\mathcal{I}_{j,k}}p_{i,j,k}\Psi_{i,j,k}+\lambda_{c}\Xi^{\text{comp}}_{j,k}\right),$
where $\Lambda_{j,k}$ represents the accuracy of the global model for task $\Upsilon_{j,k}$, and $\Xi^{\text{comp}}_{j,k}$ is the cloud aggregation cost. $\mu$ is a weight parameter that balances model accuracy and total payment.
@@
Attack Detection Tran et al. (tran2023efficient) introduce a privacy-preserving FL approach specifically designed to detect false data injection attacks (FDIAs) in smart grid systems. This method focuses on enhancing data privacy and security in cross-silo FL scenarios within the SG environment. By improving the detection efficiency and effectiveness of FDIAs, their approach ensures that data confidentiality is maintained across various entities involved in the grid, making it a robust solution for securing SG infrastructure. Zhao et al. (zhao2024federated) propose a FL algorithm that integrates client sampling and gradient projection techniques to enhance SG applications. This approach effectively addresses privacy concerns by enabling model training using fragmented data from devices spread across the grid. Furthermore, the inclusion of client sampling and gradient projection techniques leads to improved FL performance and optimization, providing a more reliable framework for handling SG data without compromising security.
@@
Anomaly Detection Kesici et al. (kesici2024detection) propose a data-driven time series anomaly detection method that leverages time-aware shapelets, a shapelet evolution graph, and segment embeddings learned through the DeepWalk algorithm. Their method demonstrates a significant improvement in detecting FDIA, with a 20-40% increase in accuracy compared to other unsupervised learning techniques. Additionally, when applied with BO-XGBoost, the model achieves a 5% higher accuracy than unmodified XGBoost, adding further robustness to the detection process. Moreover, this approach imparts physical significance to the dynamic evolution of time series models, providing valuable insights into the underlying data. Jithish et al. (jithish2023distributed) present a FL-based approach for distributed anomaly detection in SGs. Their solution enhances privacy and security by allowing for collaborative detection of anomalies without the need to centralize sensitive data. This distributed approach not only safeguards data privacy but also improves the overall stability and reliability of SG systems. By enabling effective anomaly detection across the grid, their method ensures that potential issues are identified early, contributing to more resilient SG operations.
Security Assessment Li et al. (li2024privacy) introduce a quantum-secured FL system, known as QFDSA, for dynamic security assessment in SGs. This system integrates quantum-resistant cryptography with FL to ensure both data privacy and security, particularly in the context of SG dynamic security assessments (DSA). Their approach addresses the limitations of traditional DSA methods, which may be vulnerable to privacy breaches and quantum computing attacks, thereby offering a more secure alternative for assessing grid security. Ren et al. (ren2023qfdsa) further develop QFDSA, a hybrid system that combines quantum security and FL for dynamic security assessment in SGs. By enhancing privacy and security measures, their approach tackles key challenges related to data protection and collaborative analysis in SG environments. Specifically, in Quantum Key Distribution (QKD) for FL, secure keys are exchanged between participants, typically denoted as Alice and Bob, using quantum states to ensure security. In a Measurement-Device-Independent QKD (MDI-QKD) protocol, both Alice and Bob randomly choose a measurement basis from either $K$ or $Z$ and send weak coherent-state pulses (WCPs) with intensities sampled from predefined sets. These pulses are transmitted to a trusted or untrusted relay (Charlie) which performs a Bell state measurement (BSM). The key rate $R$ for secure communication is calculated by:
(6) $R=P_{s}^{A}P_{s}^{B}\left[s_{A}e^{-s_{A}}\cdot s_{B}e^{-s_{B}}\cdot Y_{11}^{K,L}\left(1-h(e_{11}^{K,U})\right)-feQ^{Z}h(E^{Z})\right].$
Here, $P_{s}^{A}$ and $P_{s}^{B}$ resent the probabilities that Alice and Bob send WCPs with the intensity of the signal state. $Y_{11}^{K,L}$ and $e_{11}^{K,U}$ are the yield and error rates in the decoy basis $K$, respectively. $f$ is the error correction efficiency, and $Q^{Z}$ and $E^{Z}$ are the gain and error rates in the signal basis $Z$, ensuring quantum-secured keys for FL.
Ren et al. (ren2023efeddsa) propose EFedDSA, an efficient differential privacy (DP)-based horizontal FL approach for SG dynamic security assessment. This system enhances privacy protection in data-driven security assessments, addressing the challenge of data isolation and privacy concerns in traditional machine learning approaches. EFedDSA improves the efficiency and effectiveness of dynamic security assessment while maintaining data confidentiality across multiple participants in the SG infrastructure, offering a comprehensive solution for modern SG security needs.
3.3. Consumption Stage
@@
In the area of optimal power usage (rajesh2024federated; abdulla2024smart), FL facilitates the extraction of typical electricity consumption patterns from smart meter data, enabling more efficient grid management. For load forecasting (badr2023privacy; liu2023fedforecast; fekri2023asynchronous), FL-based methods have been developed to enhance the accuracy of short-term residential load predictions by utilizing distributed data from multiple sources. In the case of electricity theft detection (li2024fuse; alshehri2024deep; zafar2023step), frameworks like FedDetect employ federated learning to effectively identify fraudulent activities while preserving consumer privacy.
@@
Optimal Power Usage Rajesh et al. (rajesh2024federated) propose a FL framework designed to enhance privacy while optimizing power usage in SGs. The framework allows for collaborative learning across multiple participants without compromising individual user data. By securing power traces and enabling personalized recommendations for power usage optimization, the approach significantly improves privacy preservation in SG data analytics. This personalized aspect ensures that recommendations are tailored to the specific needs of users while maintaining data confidentiality. Similarly, Abdulla et al. (abdulla2024smart) present an adaptive FL approach for energy consumption forecasting using smart meter data. Their method improves forecasting accuracy by utilizing smart meter data without sharing raw information, ensuring user privacy. The adaptive nature of the approach allows it to adjust to individual consumption patterns, providing personalized forecasting models for each participant. This enhances both the accuracy of energy predictions and the privacy of users in the SG system.
@@
Load Forecasting Badr et al. (badr2023privacy) introduce a privacy-preserving and communication-efficient energy prediction scheme based on FL for SGs. Their approach enables accurate load forecasting while maintaining data confidentiality, which is crucial for infrastructure planning and power dispatching. By enhancing communication efficiency in the FL process, this method helps reduce power outages and equipment failures, making it an important tool for SG operations. Liu et al. (liu2023fedforecast) propose FedForecast, a FL framework designed for short-term probabilistic individual load forecasting. This framework provides comprehensive information about future load uncertainties while addressing data isolation and privacy concerns inherent in traditional centralized forecasting methods. By maintaining data confidentiality across households or smart meters, FedForecast improves the accuracy and reliability of short-term load forecasts. Fekri et al. (fekri2023asynchronous) develop an asynchronous adaptive FL algorithm for distributed load forecasting with smart meter data. This approach enhances the scalability and efficiency of load forecasting in SGs by adapting asynchronously to the data, thus addressing issues of privacy, security, and communication overhead found in traditional centralized methods. By doing so, the algorithm ensures accurate load forecasting while preserving the privacy of distributed data.
@@
Energy Theft Detection Li et al. (li2024fuse) introduce FUSE, a framework that combines FL with U-shape split learning for detecting electricity theft. FUSE identifies abnormal electricity consumption behaviors while preserving the privacy of both user data and the detection model. By integrating FL with split learning, FUSE enhances data protection and model security in smart grid systems, offering a robust solution for electricity theft detection. Alshehri et al. (alshehri2024deep) develop a deep anomaly detection framework that utilizes FL to detect electricity theft, including zero-day threats, in smart power grids. This framework preserves privacy while effectively identifying new and previously unseen attacks, thereby enhancing the security and stability of SGs. By addressing the challenge of detecting sophisticated electricity theft attempts, the framework plays a critical role in maintaining grid reliability. Zafar et al. (zafar2023step) present a FL-assisted hybrid deep learning model for electricity theft detection using smart meter data. This model enhances the security and reliability of SGs, particularly in the context of Industry 5.0. By leveraging FL to keep data decentralized, the model addresses privacy concerns while effectively detecting electricity theft. This hybrid approach offers a comprehensive solution to safeguard SGs from fraudulent activities.
@@ 4. Vulnerabilities in FL-based SG Systems
In this section, we explore the vulnerabilities present in SG systems. As illustrated in Table 2, our discussion is structured around three critical stages within SGs: generation, transmission and distribution, and consumption, as well as two vulnerability aspects, as depicted in Figure 4: privacy and security. This systematic approach allows us to comprehensively demonstrate the potential vulnerabilities associated with various applications within these stages. In each subsection, In each subsection, We begin by presenting the fundamental concepts underlying each application to establish a clear understanding of their functionalities and objectives. We then examine the specific aspects that make these applications potentially vulnerable, highlighting their inherent risks.
4.1. Vulnerabilities in Generation Stage
@@
In the context of SG, generation refers to the production of electricity from various sources, including both traditional and renewable energy (tuballa2016review). It is a crucial component of the SG pipeline, serving as the starting point for the flow of electricity throughout the entire system. As generation systems become more integrated and digitalized, privacy and security becomes increasingly important (krause2021cybersecurity). Protecting generation infrastructure from attacks is essential for maintaining the reliability and integrity of the entire SG system.
4.1.1. Energy Forecasting
A key aspect of generation in SGs is the integration of energy forecasting, particularity for renewable energy sources. SGs are designed to accommodate and efficiently manage the intermittent nature of renewable energy generation, such as solar and wind power (wang2023investigation). This integration helps reduce reliance on fossil fuels and promotes a more sustainable energy future. Specifically, energy forecasting leverages advanced machine learning techniques, including deep learning models like Long Short-Term Memory (LSTM) networks, to analyze historical data and predict future energy generation patterns (malik2022integration).
Although accurate forecasting helps maintain grid stability by predicting fluctuations in renewable energy output, it also introduces new vulnerabilities, as attackers may target the prediction process. In the context of the FL framework, Shabbir et al. (shabbir2024resilience) focus on energy forecasting using FL in SG applications. The data involved typically consists of time-series data related to energy consumption and generation patterns. one primary attack method discussed is FDIA, which can take various forms, such as scaling attacks, random noise attacks, and mean-based attacks. These attacks aim to manipulate the learning process by injecting falsified data into the FL system. The impact of these attacks is often measured using metrics like Mean Absolute Percentage Error (MAPE). For instance, experiments have shown that scaling attacks can increase MAPE from 0.193% to 32.72%, while random noise attacks can elevate it from 0.183% to 129.75% as the proportion of attacked data rises from 10% to 100%. Badr et al. (badr2022privacy) propose a privacy-preserving FL approach for net-energy forecasting, addressesing the critical need for accurate energy predictions while protecting individual privacy in SG systems. To enhance privacy protection, researchers have introduced various strategies, including secure data aggregation (fereidooni2021safelearn), differential privacy, and hybrid mechanisms. These approaches aim to mitigate potential privacy leaks that may occur through attacks on shared model parameters. FedGrid (gupta2023fedgrid) is an innovative framework that leverages FL to optimize energy distribution in SG systems. It addresses two major challenges in the energy sector: accurately predicting renewable energy generation and forecasting electric load, while simultaneously maintaining data privacy and security. By utilizing FL, FedGrid enables collaborative model training across multiple parties without compromising the privacy of individual data, a crucial factor when handling sensitive information like user electricity consumption patterns and localized weather data. FedGrid’s comprehensive analysis helps SG operators develop efficient supply strategies, optimizing the distribution of energy from both renewable and non-renewable sources.
@@ 4.1.1. Energy Forecasting
Generation Shabbir et al. (shabbir2024resilience) 2024 Energy Forecasting Time Series False Data Injection Security
@@
Badr et al. (badr2022privacy) 2022 Net-Energy Forecasting Time Series Model Inversion, Membership Inference Privacy
@@
FedGrid (gupta2023fedgrid) 2023 Net-Energy Forecasting Miscellaneous Data Inference Privacy
Moayyed et al. (moayyed2022cyber) 2022 Wind Speed Forecasting Miscellaneous Data Inference Privacy
T&D Lin et al. (lin2022incentive) 2022 State Estimation Measurements False Data Injection Security
@@
Mohammadabadi et al. (mohammadabadi2023towards) 2023 Event Classification Phasor Measurement Unit Data Inference Privacy
Jithish et al. (jithish2023distributed) 2023 Anomaly Detection Miscellaneous Data Integrity Security
Liu et al. (liu2022federated) 2022 Voltage Control Miscellaneous Data Inference Privacy
Consumption Husnoo et al. (husnoo2023secure) 2023 Load Forecasting Historical Data Integrity Security
@@
Wen et al. (wen2021feddetect) 2021 Energy Theft Detection Irish Smart Meter Inference, Model Poisoning Privacy, Security
@@
Ashraf et al. (ashraf2022feddp) 2022 Energy Theft Detection Smart Meter Data Poisoning Security
@@
Zafar et al. (zafar2023step) 2023 Energy Theft Detection Irish Smart Meter Data Poisoning Security
@@
Li et al. (li2024fuse) 2024 Energy Theft Detection Smart Meter Data Poisoning, Model Inversion Privacy, Security
4.1.2. Wind Speed Forecasting
Wind speed forecasting is a crucial process in renewable energy management, involving the prediction of future wind speeds using advanced statistical and machine learning techniques (alhussein2019microgrid). This method utilizes historical wind speed data, along with other meteorological parameters, to generate accurate predictions for various time horizons, ranging from short-term (hours ahead) to long-term (days or weeks ahead) (malik2022multi). Accurate wind speed forecasting directly contributes to improved wind power prediction, enabling the seamless integration of wind energy into the SG. By anticipating wind power output, grid operators can better optimize the balance between wind energy and other power sources, resulting in more efficient utilization of renewable resources (ghofrani2014smart).
Wind speed forecasting often relies on advanced machine learning and deep learning algorithms (lin2024prediction; sun2021solar), which can be vulnerable to adversarial attacks designed to manipulate their outputs. Attackers could exploit weaknesses in these algorithms, introducing errors in wind speed predictions and leading to inaccurate power generation forecasts. Within the FL framework, Moayyed et al. (moayyed2022cyber) propose a wind power forecasting framework integrated into the FL pipeline. The system predicts future wind power generation using multiple resource, including time-series data on wind power generation and meteorological data. Additionally, the model incorporates Sentinel-1 and Sentinel-2 satellite images for wind speed and bathymetry analysis, enhancing the spatial information available for forecasting. While the FL approach inherently provides some defense against data breaches by decentralizing sensitive information, local model updates from attackers can potentially compromise the entire system (zhang2015byzantine; liu2020reflection).
4.2. Vulnerabilities in Transmission and Distribution Stages

### v3

@@ 3.1. Generation Stage
FL enables collaborative model training across different energy producers and consumers, enhancing the accuracy of energy generation forecasts and enabling dynamic resource allocation without the need to centralize sensitive operational data (Lee, 2023; Gupta et al., 2023). In the context of energy management (Abdellatif et al., 2024), FL supports the development of intelligent systems that can predict and optimize power generation patterns while keeping sensitive user data local on smart devices. This approach improves grid efficiency and facilitates more effective demand response programs.
Energy Forecasting. In the work by (Lee, 2023), an advanced algorithm for FL is proposed to tackle the challenges of non-IID data in distributed solar energy grids. This framework is specifically designed to handle the heterogeneous nature of the data across different solar energy producers, enhancing the model’s ability to converge efficiently in decentralized settings. By addressing the variations in data distribution, the framework improves the accuracy and reliability of predictions, leading to better performance in solar energy generation and more effective resource allocation.
Energy Optimization. The FedGrid framework introduced by (Gupta et al., 2023) offers a secure FL solution for optimizing energy management within smart grids. This framework emphasizes the importance of protecting privacy and enhancing data security during the optimization process. By enabling efficient collaboration between various grid components without compromising data confidentiality, FedGrid helps optimize energy distribution and usage. The approach also addresses the key challenges associated with data sharing and collaboration in smart grids, ultimately improving the overall efficiency and reliability of the system. As shown in Figure 3, energy optimization is crucial for coordinating multiple energy resources and consumption, especially when using FL techniques to protect data privacy.
Energy Management. In (Abdellatif et al., 2024), a Secure, Distributed, and Collaborative Learning (SDCL) framework is presented to enhance the operations of smart grids. This framework focuses on maintaining data privacy and security while facilitating collaborative learning among grid components. By incorporating advanced AI and blockchain technologies, SDCL aims to optimize grid performance, improve system efficiency, and ensure reliable energy management. Additionally, (Lin and Ciou, 2024) introduces a privacy-preserving energy management framework that utilizes vertical FL for smart data cleaning in smart homes. This approach enhances user privacy while enabling effective analysis and management of smart meter data, ensuring that energy usage is optimized without compromising sensitive information.
@@ 3.2. Transmission and Distribution Stages
For energy distribution (Mourtzis et al., 2023), FL enables optimal energy distribution solutions and flexibility predictions, significantly improving the efficiency of power distribution in SGs. In the area of attack (Tran et al., 2023; Zhao et al., 2024) and anomaly detection (Kesici et al., 2024; Jithish et al., 2023), FL has demonstrated its effectiveness in identifying cyber attacks and anomalies within SG systems. This approach allows multiple stakeholders to collaboratively train robust detection models while maintaining data privacy, as raw consumption data remains local. In the realm of assessment (Li et al., 2024b; Ren et al., 2023b, a), FL has been applied to dynamic security assessment, enabling participants to collaboratively build models while preserving data privacy.
Energy Distribution. Mourtzis et al. (Mourtzis et al., 2023) introduce an FL framework designed to enhance the efficiency and security of energy distribution within smart grids, while concurrently addressing privacy concerns by enabling collaborative model training without the exchange of raw data. A key feature of their proposed system is an incentive mechanism where service providers compensate edge devices based on their contributions to designated tasks.
This mechanism is structured around two primary payoff functions. For an edge device, the payoff function balances the revenue earned—influenced by factors such as the timeliness or ”Age of Local Model” contributed—against its operational costs, which encompass local data processing, computation, and communication expenses. For a service provider, the corresponding payoff function weighs the accuracy achieved by the aggregated global model against the total payments made to participating edge devices and the costs incurred for cloud-based aggregation. The objective of this incentive structure is to encourage robust participation and high-quality contributions from edge devices while allowing service providers to obtain accurate models in a cost-effective manner.
@@
Attack Detection. Tran et al. (Tran et al., 2023) introduce a privacy-preserving FL approach specifically designed to detect false data injection attacks (FDIAs) in smart grid systems. This method focuses on enhancing data privacy and security in cross-silo FL scenarios within the SG environment. By improving the detection efficiency and effectiveness of FDIAs, their approach ensures that data confidentiality is maintained across various entities involved in the grid, making it a robust solution for securing SG infrastructure. Zhao et al. (Zhao et al., 2024) propose a FL algorithm that integrates client sampling and gradient projection techniques to enhance SG applications. This approach effectively addresses privacy concerns by enabling model training using fragmented data from devices spread across the grid. Furthermore, the inclusion of client sampling and gradient projection techniques leads to improved FL performance and optimization, providing a more reliable framework for handling SG data without compromising security.
@@
Anomaly Detection. Kesici et al. (Kesici et al., 2024) propose a data-driven time series anomaly detection method that leverages time-aware shapelets, a shapelet evolution graph, and segment embeddings learned through the DeepWalk algorithm. Their method demonstrates a significant improvement in detecting FDIA, with a 20-40% increase in accuracy compared to other unsupervised learning techniques. Additionally, when applied with BO-XGBoost, the model achieves a 5% higher accuracy than unmodified XGBoost, adding further robustness to the detection process. Moreover, this approach imparts physical significance to the dynamic evolution of time series models, providing valuable insights into the underlying data. Jithish et al. (Jithish et al., 2023) present an FL-based approach for distributed anomaly detection in SGs. Their solution enhances privacy and security by allowing for collaborative detection of anomalies without the need to centralize sensitive data. This distributed approach not only safeguards data privacy but also improves the overall stability and reliability of SG systems. By enabling effective anomaly detection across the grid, their method ensures that potential issues are identified early, contributing to more resilient SG operations.
Security Assessment. Several approaches aim to enhance security and privacy in FL for SG dynamic security assessment (DSA). Li et al. (Li et al., 2024b) introduced QFDSA, a system that integrates FL with quantum-resistant cryptography to bolster data privacy and security against advanced threats, including those from quantum computing. Building on this, Ren et al. (Ren et al., 2023b) further developed QFDSA, emphasizing the use of Quantum Key Distribution (QKD) protocols, such as Measurement-Device-Independent QKD (MDI-QKD). In these QKD systems, participants (like “Alice” and “Bob”) exchange quantum states, often via a relay ( “Charlie”), to establish highly secure cryptographic keys. The effectiveness and security of these keys, crucial for protecting the FL process, are determined by analyzing factors like signal transmission probabilities and error rates, without needing to delve into the complex underlying mathematical formulas for a general understanding.
In a distinct but related effort, Ren et al. (Ren et al., 2023a) proposed EFedDSA, an efficient horizontal FL approach that incorporates differential privacy for SG dynamic security assessment. This method strengthens privacy by protecting individual data contributions during the collaborative security analysis, thereby addressing data isolation challenges and improving the overall efficiency and confidentiality of DSA in SG environments.
3.3. Consumption Stage.
@@
In the area of optimal power usage (Rajesh et al., 2024; Abdulla et al., 2024), FL facilitates the extraction of typical electricity consumption patterns from smart meter data, enabling more efficient grid management. For load forecasting (Badr et al., 2023; Liu et al., 2023; Fekri et al., 2023), FL-based methods have been developed to enhance the accuracy of short-term residential load predictions by utilizing distributed data from multiple sources. In the case of electricity theft detection (Li et al., 2024a; Alshehri et al., 2024; Zafar et al., 2023), frameworks like FedDetect employ FL to effectively identify fraudulent activities while preserving consumer privacy.
@@
Optimal Power Usage. Rajesh et al. (Rajesh et al., 2024) propose a FL framework designed to enhance privacy while optimizing power usage in SGs. The framework allows for collaborative learning across multiple participants without compromising individual user data. By securing power traces and enabling personalized recommendations for power usage optimization, the approach significantly improves privacy preservation in SG data analytics. This personalized aspect ensures that recommendations are tailored to the specific needs of users while maintaining data confidentiality. Similarly, Abdulla et al. (Abdulla et al., 2024) present an adaptive FL approach for energy consumption forecasting using smart meter data. Their method improves forecasting accuracy by utilizing smart meter data without sharing raw information, ensuring user privacy. The adaptive nature of the approach allows it to adjust to individual consumption patterns, providing personalized forecasting models for each participant. This enhances both the accuracy of energy predictions and the privacy of users in the SG system.
@@
Load Forecasting. Badr et al. (Badr et al., 2023) introduce a privacy-preserving and communication-efficient energy prediction scheme based on FL for SGs. Their approach enables accurate load forecasting while maintaining data confidentiality, which is crucial for infrastructure planning and power dispatching. By enhancing communication efficiency in the FL process, this method helps reduce power outages and equipment failures, making it an important tool for SG operations. Liu et al. (Liu et al., 2023) propose FedForecast, a FL framework designed for short-term probabilistic individual load forecasting. This framework provides comprehensive information about future load uncertainties while addressing data isolation and privacy concerns inherent in traditional centralized forecasting methods. By maintaining data confidentiality across households or smart meters, FedForecast improves the accuracy and reliability of short-term load forecasts. Fekri et al. (Fekri et al., 2023) develop an asynchronous adaptive FL algorithm for distributed load forecasting with smart meter data. This approach enhances the scalability and efficiency of load forecasting in SGs by adapting asynchronously to the data, thus addressing issues of privacy, security, and communication overhead found in traditional centralized methods. By doing so, the algorithm ensures accurate load forecasting while preserving the privacy of distributed data.
@@
Energy Theft Detection. Li et al. (Li et al., 2024a) introduce FUSE, a framework that combines FL with U-shape split learning for detecting electricity theft. FUSE identifies abnormal electricity consumption behaviors while preserving the privacy of both user data and the detection model. By integrating FL with split learning, FUSE enhances data protection and model security in smart grid systems, offering a robust solution for electricity theft detection. Alshehri et al. (Alshehri et al., 2024) develop a deep anomaly detection framework that utilizes FL to detect electricity theft, including zero-day threats, in smart power grids. This framework preserves privacy while effectively identifying new and previously unseen attacks, thereby enhancing the security and stability of SGs. By addressing the challenge of detecting sophisticated electricity theft attempts, the framework plays a critical role in maintaining grid reliability. Zafar et al. (Zafar et al., 2023) present a FL-assisted hybrid deep learning model for electricity theft detection using smart meter data. This model enhances the security and reliability of SGs, particularly in the context of Industry 5.0. By leveraging FL to keep data decentralized, the model addresses privacy concerns while effectively detecting electricity theft. This hybrid approach offers a comprehensive solution to safeguard SGs from fraudulent activities.
@@ 4. Vulnerabilities in FL-based SG Systems
In this section, we explore the vulnerabilities present in SG systems. As illustrated in Table 2, our discussion is structured around three critical stages within SGs: generation, transmission and distribution, and consumption, as well as two vulnerability aspects, as depicted in Figure 4: privacy and security. This systematic approach allows us to comprehensively demonstrate the potential vulnerabilities associated with various applications within these stages. In each subsection, we begin by presenting the fundamental concepts underlying each application to establish a clear understanding of their functionalities and objectives. We then examine the specific aspects that make these applications potentially vulnerable, highlighting their inherent risks.
4.1. Traditional Adversarial Attack and Defense Methods
4.1.1. Vulnerabilities in Generation Stage
@@
In the context of SG, generation refers to the production of electricity from various sources, including both traditional and renewable energy (Tuballa and Abundo, 2016). It is a crucial component of the SG pipeline, serving as the starting point for the flow of electricity throughout the entire system. As generation systems become more integrated and digitalized, privacy and security become increasingly important (Krause et al., 2021). Protecting generation infrastructure from attacks is essential for maintaining the reliability and integrity of the entire SG system.
Energy Forecasting. Accurate energy forecasting is vital for integrating intermittent renewable energy sources into SGs (Wang et al., 2023), often leveraging machine learning techniques on historical and meteorological data (Malik and Pouttu, 2022). However, this reliance on data-driven models, whether centralized or distributed, introduces vulnerabilities to adversarial manipulation. FL approaches for energy forecasting attempt to address these issues, often by adapting defenses against threats similar to those seen in traditional systems. For instance, Shabbir et al. (Shabbir et al., 2024) explore how False Data Injection Attacks (FDIAs)—a common concern in centralized grid monitoring—can impact FL-based energy forecasting. Their work shows that various FDIA manifestations, such as scaling or random noise attacks on local data, can significantly degrade the FL model’s predictive accuracy. Addressing fundamental data privacy concerns, which are critical in any data-handling system, Badr et al. (Badr et al., 2022) propose a privacy-preserving FL approach for net-energy forecasting. They employ techniques like secure data aggregation and differential privacy to protect against inference attacks on shared model parameters; these defense strategies are also foundational in protecting sensitive data in many centralized machine learning applications. Similarly, the FedGrid framework (Gupta et al., 2023) uses FL for renewable energy and load forecasting, inherently aiming to mitigate the privacy risks associated with traditional centralized data aggregation, where pooling sensitive user consumption and weather data would make it a prime target for conventional data breaches. FedGrid enables collaborative training without direct raw data sharing, thus addressing a key vulnerability of centralized architectures.
@@
Shabbir et al. (Shabbir et al., 2024) 2024 Energy Forecasting Time Series False Data Injection Security
@@
Badr et al. (Badr et al., 2022) 2022 Net-Energy Forecasting Time Series Model Inversion, Membership Inference Privacy
@@
FedGrid (Gupta et al., 2023) 2023 Net-Energy Forecasting Miscellaneous Data Inference Privacy
T&D
Lin et al. (Lin et al., 2022) 2022 State Estimation Measurements False Data Injection Security
@@
Mohammadabadi et al. (Mohammadabadi et al., 2023) 2023 Event Classification Phasor Measurement Unit Data Inference Privacy
Jithish et al. (Jithish et al., 2023) 2023 Anomaly Detection Miscellaneous Data Integrity Security
Liu et al. (Liu and Wu, 2022) 2022 Voltage Control Miscellaneous Data Inference Privacy
Consumption
Husnoo et al. (Husnoo et al., 2023) 2023 Load Forecasting Historical Data Integrity Security
@@
Wen et al. (Wen et al., 2021) 2021 Energy Theft Detection Irish Smart Meter Inference, Model Poisoning Privacy, Security
@@
Ashraf et al. (Ashraf et al., 2022) 2022 Energy Theft Detection Smart Meter Data Poisoning Security
@@
Zafar et al. (Zafar et al., 2023) 2023 Energy Theft Detection Irish Smart Meter Data Poisoning Security
@@
Li et al. (Li et al., 2024a) 2024 Energy Theft Detection Smart Meter Data Poisoning, Model Inversion Privacy, Security
4.1.2. Vulnerabilities in Transmission and Distribution Stages